OpenStudy (anonymous):

identify slope and y intercept of the following equation 6x-5y=-20

OpenStudy (anonymous):

slope = 6/5, y intercept = 4

OpenStudy (anonymous):

express the following equation in slope intercept form 7x-5y=-35